Shift report 20011019 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Oct 19 2001 - 00:09:23 EDT

  • Next message: Apache: "Shift report 20011019 00:00-08:00"

    ------------------------------------------------------------
    Date: Fri Oct 19 00:09:23 2001, Shift: 16:00-24:00
    Supervisor: DO, with: CEJ 
    Log book#10 pages 29 - 31
    
    Shift Summary:
    ------------------------------------------------------------
    1. Machine status:
       Nice store of 1h15 or so from 20:40 to 21:55   
       ZDC rates: start ~ 840, end ~ 450
       Beam Intensities: Blue 22.3x10^9, Yellow 19.7x10^9 
    
    2. Achievements:
       Data runs       Trig1 - Trig4 - Trig5 - Trig6
         run 5431      16853   16995   16614    5465 
         run 5432     259654  263427  249980  144736  
    
       (Trigger6 is vertex AND mult)
    		
       Pedestal run  5430
       TPC pedestal  5433
       DVM run       5434
    
    3. Detector problems:
       No problems except an IO server problem at beginning 
       of the shift (trigger 7 was not triggering). FV fixed 
       it directly in the FEH: rebooted the VME processor.
       DCs anode voltages tripped at end of run 5432 (beams
       lost without any warnings).
       D3 and D5 cathodes tripped too. Turned them off and on
       with BRAHVO, ok now.
         
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Fri Oct 19 2001 - 14:05:33 EDT